QUESTION 1

Test marks for ten students are 2, 3, 6, 7, 7, x, 5, 5, 8, 9

(i) Given that the average mark is 6, find the value of x.

(ii) Calculate the standard deviation for these set of marks

QUESTION 2

Solve the following pair of equations by Cramer’s rule.

1) 3x + 5y = 2
2) 2x-7y = - 1
1
Expert's answer
2019-08-11T09:54:10-0400

1.i)


"x+2+3+6+7+7+5+5+8+9=6(10)"

"x+52=60"

"x=8"

ii)


"(2-6)^2+(3-6)^2+2(5-6)^2+(6-6)^2+2(7-6)^2+2(8-6)^2+(9-6)^2=46"

The standard deviation for these set of marks:


"s=\sqrt{\frac{46}{9-1}}=2.4"

 2. The coefficient matrix is


"\begin{bmatrix}
 3 & 5 \\
 2 & -7
\end{bmatrix}"

And


"\begin{vmatrix}
 3 & 5 \\
 2 & -7
\end{vmatrix}=3(-7)-2(5)=-31"

So,


"x=\frac{\begin{vmatrix}
 2 & 5 \\
 -1 & -7
\end{vmatrix}}{-31}=\frac{(2)(-7)-(-1)(5)}{-31}=\frac{9}{31}"


"y=\frac{\begin{vmatrix}
 3 & 2 \\
 2 & -1
\end{vmatrix}}{-31}=\frac{(3)(-1)-(2)(2)}{-31}=\frac{7}{31}"